40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Dec 10 10:33:46 2010, Zach, Laser, GYRO, update 
    Reply  Fri Dec 10 13:43:07 2010, Koji, Laser, GYRO, update 
Message ID: 1203     Entry time: Fri Dec 10 10:33:46 2010     Reply to this: 1208
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: update 

 In light of Koji's work, I took new OLTFs of both loops last night. They look almost identical to the old ones, as we should expect if the loops are just below the gain instability point. The gain knob settings on the PDH boxes are:

  • GCCW = 0.00
  • GCW = 3.80

Before the measurements, I adjusted the input power to ensure that we had the maximum optical gain with the CCW servo at 0.00.

Also, when turning the slow loop back on after doing the EOM measurements, I noticed that there was a (slow) error signal offset of about 4000 cts. Koji had recently put in a manual offset of the same amount to counteract what was an offset in the other direction before. I turned this off and was able to get the error signal back to around 0 by clearing the servo memory (i.e. just turning the filter block off and on and then reengaging the output).

I am working on the low-frequency part of the noise budget plot, for which the simple code is done, but it appears that the DAQ is busted again (see next entry). For this, the NB code will pull a series of second-trend data over a longer period, and the low-frequency spectrum produced with it will be stitched to the other one we have been generating.

ELOG V3.1.3-